# centertableaux,boxsize=.5em

### v1

@@ 3.4.1 $A_{5}/\mathbb{Z}_{2}$ symmetry
Consider an alternating group $A_{5}$ and its subgroup $K=\mathbb{Z}_{2}$. Since $A_{5}$ is simple, $K$ is not a normal subgroup and $A_{5}/\mathbb{Z}_{2}$ is non-invertible. This is a minimal coset symmetry. Also, $A_{5}$ does not have a subgroup $H$ with order $30=|A_{5}|/|\mathbb{Z}_{2}|$, so $A_{5}$ cannot be expressed as the bicrossed product $H\Join K$ with $K=\mathbb{Z}_{2}$. So $A_{5}/\mathbb{Z}_{2}$ is a candidate of the coset non-invertible symmetry that does not admit realization in SPT phases.
@@
The $A_{5}$ gauge theory has the anyons labeled by $([g],\pi)$ with $[g]\in\text{Conj}(G)$ is the conjugacy class of $g\in G$, and $\pi$ is an irrep of the centralizer $Z(g)$ of $g\in G$. The anyons with $g=1$ corresponds to the electric particles carrying the irrep of $G$.
@@ 3.5 Dynamical scenario: symmetric gapped phases
In Sec. 3.2 we have seen that the coset symmetry $G/K$ with the twist $\omega_{D+1}$ admits realization in SPT phases, if $G$ is expressed as a bicrossed product $G=H\Join K$ with a subgroup $H\subset G$ such that $[\omega_{D+1}|_{H}]=0$. In this section, we will show that once a finite group $G$ is expressed as a bicrossed product $G=H\Join K$, the $G/K$ symmetry in $D\geq 3$ dimensions with any twist $\omega_{D+1}$ always admits realization in symmetry-preserving gapped phases.
To see this, we again consider the symmetry TQFT where the bulk TQFT is given by the $G$ gauge theory with twist $\omega_{D+1}$. The symmetry-preserving gapped phase corresponds to the thin interval of the bulk gauge theory sandwich by a pair of gapped boundaries: one boundary breaks the gauge group $G\to K$, and the other breaks $G\to H$. Since $[\omega_{D+1}|_{H}]$ can still be nontrivial in general, the $G\to H$ boundary is realized by the $D$-dimensional non-invertible TQFT with $H$ symmetry carrying the anomaly $[\omega_{D+1}|_{H}]$. Such an anomalous TQFT can be generally obtained by a finite gauge theory through the group extension of $H$ by a finite gauge group [52, 53]. By shrinking the size of the interval, we get a symmetry-preserving gapped phase with $(H\Join K)/K$ symmetry with generic twist.
@@ 3.6 Dynamical scenario: symmetry-enforced gaplessness
3.6.1 Symmetry-enforced gaplessness of finite coset symmetry
In the previous section, we have seen that finite $(H\Join K)/K$ symmetry with generic twist can be realized in symmetry-preserving gapped phases. Here we show the inverse statement; if a finite group $G$ does not admit the expression in terms of bicrossed product $G=H\Join K$ with any subgroup $H\subset G$, the system preserving the coset symmetry $G/K$ must be gapless.
As an example, let us explicitly see that the $A_{5}/\mathbb{Z}_{2}$ symmetry exhibits symmetry-enforced gaplessness. Suppose that there exists a gapped phase in $D$ spacetime dimensions with $A_{5}/\mathbb{Z}_{2}$ symmetry, where the $(D-2)$-form symmetry generated by the $\mathbb{Z}_{2}$ Wilson line is unbroken. We can gauge the $(D-2)$-form symmetry generated by the Wilson line, which brings the theory back to the gapped phase with $A_{5}$ symmetry. Then the dual $\mathbb{Z}_{2}=\{1,s\}$ symmetry must be broken. So the $A_{5}$ symmetry must have a nontrivial SSB pattern to the subgroup $A_{5}\to H$. Now there is no subgroup of $A_{5}$ with order $|A_{5}|/|\mathbb{Z}_{2}|=30$, so there exists $g\in A_{5}$ with $g\neq s$ which is broken. This implies that the 0-form coset non-invertible symmetry $\tilde{U}_{g}$ of the original theory is broken. This shows that the gapped phase with $A_{5}/\mathbb{Z}_{2}$ must spontaneously break either the $(D-2)$-form $\mathbb{Z}_{2}$ symmetry or the 0-form non-invertible symmetry, and thus leads to the symmetry-enforced gaplessness.
@@ 3.6.2 Symmetry-enforced gaplessness of continuous coset symmetry
For example, take $D=4$ and $G=SU(2)$ with $\omega_{5}$ given by Witten anomaly [57],101010 Witten anomaly is a fermionic anomaly which is not described by group cohomology, but rather expressed as the topological term $\eta\cup p_{1}(SU(2))$ for spin structure $\eta$ and $p_{1}(SU(2))$ is the first Pontryagin class of $SU(2)$ bundles describing the instanton number of $SU(2)$. Hence, the coset symmetry presented here involves the twist by the spin invertible phase, which is a fermionic generalization of the twisted coset symmetry $(G,K,\omega_{D+1},\alpha_{D})$ discussed in this paper. The topological term for the inflow of Witten anomaly corresponds to the Gu-Wen supercohomology SPT phase in (4+1)D, characterized by the fourth cohomology class $H^{4}(BSU(2),\mathbb{Z}_{2})=\mathbb{Z}_{2}$ generated by the mod 2 reduction of the Pontryagin class $p_{1}(SU(2))$ [58]. and $K$ given by any finite subgroup of $SU(2)$. Such $K$ is non-anomalous since finite group symmetry does not have nontrivial instanton number on $S^{4}$, while Witten anomaly means that there is an odd number of fermion zero modes in the presence of background with minimal $SU(2)$ instanton number on $S^{4}$ [57]. Suppose there were any gapped system with such anomalous coset symmetry, then by gauging $\text{Rep}(K)$ we would recover a gapped system with anomalous $G$ symmetry. However, such system would be forbidden by Witten anomaly, since a fermion parity $(-1)^{F}$ transformation changes the sign of the partition function on $S^{4}$ with $SU(2)$ instanton background and thus the partition function vanishes, which cannot happen for TQFT partition functions on $S^{4}$. Therefore we conclude that any system with such anomalous coset symmetry must be gapless.
@@ 3.7 Dynamical scenario: gapped phases with broken symmetries
3.7.1 Gapped phases with unbroken 0-form symmetry but broken $(D-2)$-form symmetry
Let us also describe another dynamical scenario where the system is gapped and symmetric under 0-form symmetry, but the $(D-2)$-form symmetry is spontaneously broken. In $D\geq 3$, any finite coset symmetry can be realized in gapped phases under such symmetry breaking.
@@ 3.7.2 Gapped phases with broken 0-form symmetries
#$\#\text{vacua}\in\frac{|G|}{|K|}\mathbb{Z}~{}.$ (3.4)
@@
As an example, consider spontaneously broken coset symmetry $(G=S_{3}=\mathbb{Z}_{3}\rtimes\mathbb{Z}_{2},K=\mathbb{Z}_{2},\omega_{D+1}=0,\alpha_{D}=0)$. For the coset symmetry to be broken, we choose the reference boundary to be the $e$-condensed boundary, where the Wilson lines are the irreducible representations $1$, $1$-dimensional $\mathbb{Z}_{2}$ Wilson line $W$, and the $2$-dimensional representation $\sigma$ that combines the charge-(±$(\pm 1)$ representations of $\mathbb{Z}_{3}$. The Wilson lines that can extend between the two boundaries are those that decompose under $\mathbb{Z}_{2}$ contains the identity, and they are $1,\sigma$. Thus there are $1+2=3$ point operators (counting with dimension), which give rise to $3$ vacua, in agreement with (3.4).
@@ 4.1 Non-anomalous case: $K$ gauge theory with $(H\Join K)/K$ symmetry
Since $G=H\Join K$ is a bicrossed product, each state |$\ket{g}$ with $g\in G$ can be uniquely expressed by a pair |$\ket{g}=\ket{h,k}$, $h\in H,k\in K,g=hk$. The theory after gauging $K$ is given by
@@
|$\ket{\text{GS}}=\sum_{\{g_{v}\}}\bigotimes_{v}\ket{g_{v}}\bigotimes_{e=\langle vv^{\prime}\rangle}\ket{k_{v}k^{-1}_{v^{\prime}}}$ (4.11)
@@
Figure 5: The edges nearby a vertex $v$ and a plaquette $p$.
@@ 4.1.1 SPT phases with $(H\Join K)/K$ symmetry
The Hamiltonian (4.10) gives an example of SPT phases with the coset $(H\Join K)/K$ symmetry. The other SPT phases with the same coset symmetry can be obtained by using the disentanglers of the $H$ SPT phases. To see this, let us consider the trivial $H$ SPT phase on a lattice with the local Hilbert space {|$\{\ket{h},h\in H\}$ on each vertex,
@@ 4.2 Anomalous case: $\mathbb{Z}_{2}$ gauge theory with $A_{5}/\mathbb{Z}_{2}$ symmetry
Since $sQs=Q$, the gauged Hamiltonian $H_{A_{5}/\mathbb{Z}_{2}}$ commutes with the $\mathbb{Z}_{2}$ Gauss law. This Hamiltonian has the coset $A_{5}/\mathbb{Z}_{2}$ symmetry
@@ 5.2 Future directions
In general spacetime dimension $D$, the coset symmetry describes a fusion $(D-1)$-category. Higher fusion categories are mysterious and it would be interesting to use the coset symmetry as specified by the data $(G,K,\omega_{D+1},\alpha_{D})$ to explore properties of higher fusion categories such as their consistency conditions for fusion and braiding. For example, consistency condition on fusion can from the cocycle conditions for $\omega_{D+1}$.
@@ Acknowledgment
We thank Yichul Choi, Sakura Schafer-Nameki and Matthew Yu for discussions. P.-S.H. is supported by Department of Mathematics King’s College London. R.K. is supported by the U.S. Department of Energy through grant number DE-SC0009988 and the Sivian Fund. C.Z. is supported by the Harvard Society of Fellows and the Simons Collaboration on Ultra Quantum Matter.
@@ Appendix A Anomaly-Free Condition for $(H\Join K)/K$ Coset Symmetries
Suppose there exists such irrep $\rho$ acting on the vector space $V$, which decomposes into $\rho=1\oplus\dots$ under $H,K$. Since the decomposition under $K$ contains a trivial representation, one can take a state |$\ket{\psi}$ satisfying $\rho(k)\ket{\psi}=\ket{\psi}$ for any $k\in K$.
Then, let us take a vector space $\tilde{V}\subset V$ spanned by $\rho(h)\ket{\psi}$ with any $h\in H$. One can immediately check that irrep $\rho$ acts within $\tilde{V}$. Pick some state $\rho(h^{\prime})\ket{\psi}\in\tilde{V}$ with $h^{\prime}\in H$. For $g=hk\in G,h\in H,k\in K$,
$\rho(g)\cdot\rho(h^{\prime})\ket{\psi}=\rho(h)\rho(k)\rho(h^{\prime})\ket{\psi}=\rho(hh^{\prime\prime})\ket{\psi}\in\tilde{V},$ (A.1)
with $kh^{\prime}=h^{\prime\prime}k^{\prime\prime}$ for some $h^{\prime\prime}\in H,k^{\prime\prime}\in K$. Since $\rho$ is an irrep, we get $V=\tilde{V}$. This implies any state of $V$ is expressed as a linear combination of states in the form of $\rho(h)\ket{\psi}$.
Now since the decomposition of $\rho$ contains a trivial representation under $H$, there exists a non-vanishing state |$\ket{\lambda}=\sum_{h\in H}\lambda_{h}\rho(h)\ket{\psi}$ satisfying $\rho(h)\ket{\lambda}=\ket{\lambda}$ for any $h\in H$. This implies
|$\ket{\lambda}=\frac{1}{|H|}\sum_{h\in H}\rho(h)\ket{\lambda}=\left(\frac{1}{|H|}\sum_{h^{\prime}\in H}\lambda_{h^{\prime}}\right)\left(\sum_{h\in H}\rho(h)\right)\ket{\psi}\propto\left(\sum_{h\in H}\rho(h)\right)\ket{\psi}$ (A.2)
So, one can take $\lambda_{h}=1$ for any $h\in H$ for the definition of |$\ket{\lambda}$. One can check that |$\ket{\lambda}$ is invariant under the action of $k\in K$. We have
$\rho(k)\ket{\lambda}=\sum_{h^{\prime}\in H}\rho(kh^{\prime})\ket{\psi}$ (A.3)
One can immediately see that $kh_{1}$ and $kh_{2}$ with $h_{1}\neq h_{2}$ leads to the expression $kh_{i}=h^{\prime}_{i}k^{\prime}_{i}$ with distinct elements $h^{\prime}_{1}\neq h^{\prime}_{2}$ of $H$ [^12] Let us check the contraposition. Suppose that $kh_{1}=hk_{1},kh_{2}=hk_{2}$ with $h\in H$. This implies $k_{1}h_{1}=k_{2}h_{2}$, so $h_{1}h^{-1}_{2}=k_{1}^{-1}k_{2}$. Since $H\cap K=\{\text{id}\}$ we get $h_{1}=h_{2}$. This implies
$\rho(k)\ket{\lambda}=\sum_{h^{\prime}\in H}\rho(kh^{\prime})\ket{\psi}=\sum_{h\in H}\rho(h)\ket{\psi}=\ket{\lambda}$ (A.4)
Then, this state |$\ket{\lambda}$ realizes the 1d irrep of $G$. For $g=hk\in G$,
$\rho(g)\ket{\lambda}=\rho(h)\rho(k)\ket{\lambda}=\ket{\lambda}.$ (A.5)

### v3

@@ 3.4.1 $A_{5}/\mathbb{Z}_{2}$ symmetry
Consider an alternating group $A_{5}$ and its subgroup $K=\mathbb{Z}_{2}$. We will focus on the case with $D=3$ for simplicity. Since $A_{5}$ is simple, $K$ is not a normal subgroup and $A_{5}/\mathbb{Z}_{2}$ is non-invertible. This is a minimal coset symmetry. Also, $A_{5}$ does not have a subgroup $H$ with order $30=|A_{5}|/|\mathbb{Z}_{2}|$, so $A_{5}$ cannot be expressed as the bicrossed product $H\Join K$ with $K=\mathbb{Z}_{2}$. So $A_{5}/\mathbb{Z}_{2}$ is a candidate of the coset non-invertible symmetry that does not admit realization in SPT phases.
@@
The (2+1)D $A_{5}$ gauge theory has the anyons labeled by $([g],\pi)$ with $g]\in\text{Conj}(G)$ is the conjugacy class of $g\in G$, and $\pi$ is an irrep of the centralizer $Z(g)$ of $g\in G$. The anyons with $g=1$ corresponds to the electric particles carrying the irrep of $G$.
@@ 3.5 Dynamical scenario: symmetric gapped phases
In Sec. 3.2 we have seen that the coset symmetry $G/K$ with the twist $\omega_{D+1}$ admits realization in SPT phases, if $G$ is expressed as a bicrossed product $G=H\Join K$ with a subgroup $H\subset G$ such that $\omega_{D+1}|_{H}]=0$. In this section, we will show that once a finite group $G$ is expressed as a bicrossed product $G=H\Join K$, the $G/K$ symmetry in $D\geq 3$ dimensions with any twist $\omega_{D+1}$ always admits realization in symmetry-preserving gapped phases.
To see this, we again consider the symmetry TQFT where the bulk TQFT is given by the $G$ gauge theory with twist $\omega_{D+1}$. The symmetry-preserving gapped phase corresponds to the thin interval of the bulk gauge theory sandwich by a pair of gapped boundaries: one boundary breaks the gauge group $G\to K$, and the other breaks $G\to H$. Since $\omega_{D+1}|_{H}]$ can still be nontrivial in general, the $G\to H$ boundary is realized by the $D$-dimensional non-invertible TQFT with $H$ symmetry carrying the anomaly $\omega_{D+1}|_{H}]$. Such an anomalous TQFT can be generally obtained by a finite gauge theory through the group extension of $H$ by a finite gauge group [47, 48]. By shrinking the size of the interval, we get a symmetry-preserving gapped phase with $(H\Join K)/K$ symmetry with generic twist.
@@ 3.6 Dynamical scenario: symmetry-enforced gaplessness
In the previous section, we have seen that finite $(H\Join K)/K$ symmetry with generic twist can be realized in symmetry-preserving gapped phases. Here we study the inverse statement; if a finite group $G$ does not admit the expression in terms of bicrossed product $G=H\Join K$ with any subgroup $H\subset G$, the system preserving the coset symmetry $G/K$ must be gapless. We show this statement fully in general in spacetime dimensions $D\geq 3$.
3.6.1 Symmetry-enforced gaplessness: an example
As an example, let us explicitly see that the $A_{5}/\mathbb{Z}_{2}$ symmetry exhibits symmetry-enforced gaplessness. Suppose that there exists a gapped phase in $D$ spacetime dimensions with $A_{5}/\mathbb{Z}_{2}$ symmetry, where the (D2)(D-2)-form symmetry generated by the $\mathbb{Z}_{2}$ Wilson line is unbroken. We can gauge the (D2)(D-2)-form symmetry generated by the Wilson line, which brings the theory back to the gapped phase with $A_{5}$ symmetry. Then the dual $\mathbb{Z}_{2}=\{1,s\}$ symmetry must be broken. So the $A_{5}$ symmetry must have a nontrivial SSB pattern to the subgroup $A_{5}\to H$. Now there is no subgroup of $A_{5}$ with order $|A_{5}|/|\mathbb{Z}_{2}|=30$, so there exists $g\in A_{5}$ with $g\neq s$ where $g,gs$ are broken. This implies that the 0-form coset non-invertible symmetry $\tilde{U}_{g}$ of the original theory is broken. This shows that the gapped phase with $A_{5}/\mathbb{Z}_{2}$ must spontaneously break either the (D2)(D-2)-form $\mathbb{Z}_{2}$ symmetry or the 0-form non-invertible symmetry, and thus leads to the symmetry-enforced gaplessness.
3.6.2 General statement about symmetry-enforced gaplessness
This argument can be extended to generic finite coset symmetry $G/K$, which may or may not be twisted. The argument for spacetime dimensions $D\geq 3$ is fully general.
In the case of $D=2$, the statement is shown when any subgroup $K^{\prime}\subset K$ has trivial 2nd group cohomology, $H^{2}(BK^{\prime},U(1))=0$. For instance $K=\mathbb{Z}_{N}$ satisfies this property, so $G/\mathbb{Z}_{N}$ symmetry in (1+1)D exhibits symmetry-enforced gaplessness when $G$ cannot be expressed as $H\Join\mathbb{Z}_{N}$.
Below, we consider a gapped system with $G/K$ symmetry preserving the $\text{Rep}(K)$ Wilson lines, and assume that $G$ cannot be expressed as a bicrossed product $H\Join K$. We then show that some 0-form symmetry in $G/K$ must be broken, therefore $G/K$ symmetry cannot be fully preserved in gapped phases.
When $\text{Rep}(K)$ (D2)(D-2)-form symmetry is preserved in $D\geq 3$, we claim that gauging $\text{Rep}(K)$ symmetry always yields the dual $K$ symmetry fully broken spontaneously. To see this, suppose we obtain a phase where $K^{\prime}\subset K$ symmetry is preserved. Gauging the $K$ symmetry back leads to finite group $K^{\prime}$ gauge theory. In such theory above $D\geq 3$, the $K^{\prime}$ Wilson lines are deconfined, and the $\text{Rep}(K^{\prime})$ subcategory symmetry is spontaneously broken. This implies that $K^{\prime}=\{\text{id}\}$. The same argument applies in $D=2$ when $\text{Rep}(K)$ has no nontrivial SPTs, i.e. $H^{2}(BK^{\prime},U(1))=0$ for any subgroup $K^{\prime}\subset K$.
@@ 3.6.3 Symmetry-enforced gaplessness of continuous coset symmetry
For example, take $D=4$ and $G=SU(2)$ with $\omega_{5}$ given by Witten anomaly [52], [^7] Witten anomaly is a fermionic anomaly which is not described by group cohomology, but rather expressed as the topological term $\eta\cup p_{1}(SU(2))$ for spin structure $\eta$ and $p_{1}(SU(2))$ is the first Pontryagin class of $SU(2)$ bundles describing the instanton number of $SU(2)$. Hence, the coset symmetry presented here involves the twist by the spin invertible phase, which is a fermionic generalization of the twisted coset symmetry $(G,K,\omega_{D+1},\alpha_{D})$ discussed in this paper. The topological term for the inflow of Witten anomaly corresponds to the Gu-Wen supercohomology SPT phase in (4+1)D, characterized by the fourth cohomology class $H^{4}(BSU(2),\mathbb{Z}_{2})=\mathbb{Z}_{2}$ generated by the mod 2 reduction of the Pontryagin class $p_{1}(SU(2))$ [53]. and $K$ given by any finite subgroup of $SU(2)$. Such $K$ is non-anomalous since finite group symmetry does not have nontrivial instanton number on $S^{4}$, while Witten anomaly means that there is an odd number of fermion zero modes in the presence of background with minimal $SU(2)$ instanton number on $S^{4}$ [52]. Suppose there were any gapped system with such anomalous coset symmetry, then by gauging $\text{Rep}(K)$ we would recover a gapped system with anomalous $G$ symmetry. However, such system would be forbidden by Witten anomaly, since a fermion parity $(-1)^{F}$ transformation changes the sign of the partition function on $S^{4}$ with $SU(2)$ instanton background and thus the partition function vanishes, which cannot happen for TQFT partition functions on $S^{4}$. Therefore we conclude that any system with such anomalous coset symmetry must be gapless.
@@ 3.7 Dynamical scenario: gapped phases with broken symmetries
3.7.1 Gapped phases with unbroken 0-form symmetry but broken (D2)(D-2)-form symmetry
Let us also describe another dynamical scenario where the system is gapped and symmetric under 0-form symmetry, but the (D2)(D-2)-form symmetry is spontaneously broken. In $D\geq 3$, any finite coset symmetry can be realized in gapped phases under such symmetry breaking.
@@ 3.7.2 Gapped phases with broken 0-form symmetries
$\text{vacua}\in\frac{|G|}{|K|}\mathbb{Z}~.$ (3.4)
@@
As an example, consider spontaneously broken coset symmetry $(G=S_{3}=\mathbb{Z}_{3}\rtimes\mathbb{Z}_{2},K=\mathbb{Z}_{2},\omega_{D+1}=0,\alpha_{D}=0)$. For the coset symmetry to be broken, we choose the reference boundary to be the $e$-condensed boundary, where the Wilson lines are the irreducible representations $1$, 11-dimensional $\mathbb{Z}_{2}$ Wilson line $W$, and the $2$-dimensional representation $\sigma$ that combines the charge-$(\pm 1)$ representations of $\mathbb{Z}_{3}$. The Wilson lines that can extend between the two boundaries are those that decompose under $\mathbb{Z}_{2}$ contains the identity, and they are $1,\sigma$. Thus there are $1+2=3$ point operators (counting with dimension), which give rise to $3$ vacua, in agreement with (3.4).
@@ 4.1 Non-anomalous case: $K$ gauge theory with $(H\Join K)/K$ symmetry
[…]
@@
[…]
@@
Figure 5: The edges nearby a vertex $v$ and a plaquette pp.
@@ 4.1.1 SPT phases with $(H\Join K)/K$ symmetry
[…]
@@ 4.2 Anomalous case: $\mathbb{Z}_{2}$ gauge theory with $A_{5}/\mathbb{Z}_{2}$ symmetry
The gauged Hamiltonian $H_{A_{5}/\mathbb{Z}_{2}}$ commutes with the $\mathbb{Z}_{2}$ Gauss law. This Hamiltonian has the coset $A_{5}/\mathbb{Z}_{2}$ symmetry
@@ 5.2 Future directions
In general spacetime dimension $D$, the coset symmetry describes a fusion (D1)(D-1)-category. Higher fusion categories are mysterious and it would be interesting to use the coset symmetry as specified by the data $(G,K,\omega_{D+1},\alpha_{D})$ to explore properties of higher fusion categories such as their consistency conditions for fusion and braiding. For example, consistency condition on fusion can originate from the cocycle conditions for $\omega_{D+1}$.
@@ Acknowledgment
We thank Yichul Choi, Kansei Inamura, Sakura Schafer-Nameki and Matthew Yu for helpful discussions. P.-S.H. is supported by Department of Mathematics King’s College London. R.K. is supported by the U.S. Department of Energy through grant number DE-SC0009988 and the Sivian Fund. C.Z. is supported by the Harvard Society of Fellows and the Simons Collaboration on Ultra Quantum Matter. The authors of this paper were ordered alphabetically.
@@ Appendix A Anomaly-Free Condition for $(H\Join K)/K$ Coset Symmetries
[…]
One can immediately see that $kh_{1}$ and $kh_{2}$ with $h_{1}\neq h_{2}$ leads to the expression $kh_{i}=h^{\prime}_{i}k^{\prime}_{i}$ with distinct elements $h^{\prime}_{1}\neq h^{\prime}_{2}$ of HH [^9] Let us check the contraposition. Suppose that $kh_{1}=hk_{1},kh_{2}=hk_{2}$ with $h\in H$. This implies $k_{1}h_{1}=k_{2}h_{2}$, so $h_{1}h^{-1}_{2}=k_{1}^{-1}k_{2}$. Since $H\cap K=\{\text{id}\}$ we get $h_{1}=h_{2}$. This implies
[…]